OpenStudy (theopenstudyowl):

Giving medals to correct answer. Joe is a fish thrower who throws fish into a chute at a processing plant. Joe misses the chute 20% of the time. Using the normal approximation with the continuity correction, which of these correctly represents the probability Joe will miss fewer than 80 times if he throws 500 fish? P(X < 80) for N(100, 8.94) P(X < 79.5) for N(100, 8.94) P(X 80) for N(100, 8.94) P(X < 80.5) for B(500, .2) P(X < 80) for B(500, .2)

OpenStudy (theopenstudyowl):

@Luigi0210

Directrix (directrix):

@kropot72

OpenStudy (theopenstudyowl):

@ganeshie8

OpenStudy (kropot72):

The mean number of misses is given by n * p = 500 * 0.2=100. The standard deviation is given by: \[\large S.D.=\sqrt{np(1-p)}=\sqrt{500\times0.2\times0.8}=8.94\] To apply the continuity correction, we need to add 0.5 to 80 giving 80.5. It appears, therefore , that none of the choices is correct.